OpenStudy (anonymous):

A drug trial is testing the effectiveness of two drugs. If 30 patients are given Drug A, 10 patients are given Drug B, and 90 patients are given a placebo?

OpenStudy (anonymous):

A: 4/ 13 B: 3/13 C: 9/13 D: 1/13

OpenStudy (anonymous):

9/13 are given a placebo. 3/13 are given Drug A 1/13 are given Drug B and 4/13 aren't given a placebo. The question didn't seem too clear.

OpenStudy (anonymous):

A drug trial is testing the effectiveness of two drugs. If 30 patients are given Drug A, 10 patients are given Drug B, and 90 patients are given a placebo,what is the probability that a patient will not be given a placebo?

OpenStudy (anonymous):

Ok, so do you understand why that is the answer?

OpenStudy (anonymous):

ooohh yeah Ty

OpenStudy (anonymous):

np!

OpenStudy (anonymous):

Probability = Number of favorable outcomes/Total number of outcomes. No. of favorable outcomes = 30 + 10 = 40 Total number of outcomes = 90+10+30 = 130 => Probability = 40/1130 = 4/13. So, we arrive at option (A).